CAN SOMEONE PLEASE ASSIST:
PLEASE SOLVE:
6X+4(3X-2)=6X+2

Found 2 solutions by tutorcecilia, Scriptor:
Answer by tutorcecilia(2152)

6X+4(3X-2)=6X+2
6x+12x-8=6x+2
18x-8=6x+2
18x-8+8=6x+2+8
18x=6x+10
18x-6x=6x-6x+10
12x=10
x=10/12
x=5/6
.
Check by plugging (x=5/6) back into the original equation.

Answer by Scriptor(36)

6X+4(3X-2)=6X+2
Cancel out the 6x at both sides to obtain: 4(3x-2) = 2
4(3x-2) = 4*3x - 4*2 = 12x-8
4(3x-2) = 2
=> 12x-8 = 2 => 12x = 10 => x = 10/12